OpenStudy (anonymous):

derivative of t^(t^2/2) ?

OpenStudy (anonymous):

? \[\huge \frac{d}{dt}t^{\frac{t^2}2}\] this?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Well, start with this... \[\huge y =t^{\frac{t^2}2} \]

OpenStudy (anonymous):

And get the natural logarithm of both sides ^.^ Logs are great, they turn exponents into products... sort of... \[\huge \ln \ y =\ln \ t^{\frac{t^2}2} \]

OpenStudy (anonymous):

And there's this one neat property of logarithms that allows you to do this... \[\huge \ln \ y = \frac{t^2}{2} \ \ln t\] Now feel free to differentiate implicitly... and you can thank me later :P

OpenStudy (anonymous):